Janes Company provided the following information on intangible assets:

  1. A patent was purchased from the Lou Company for $1,050,000 on January 1, 2019. Janes estimated the remaining useful life of the patent to be 10 years. The patent was carried on Lou’s accounting records at a net book value of $420,000 when Lou sold it to Janes.
  2. During 2021, a franchise was purchased from the Rink Company for $570,000. The contractual life of the franchise is 10 years and Janes records a full year of amortization in the year of purchase.
  3. Janes incurred research and development costs in 2021 as follows:
Materials and supplies $ 147,000
Personnel 187,000
Indirect costs 67,000
Total $ 401,000
  1. Effective January 1, 2021, based on new events that have occurred, Janes estimates that the remaining life of the patent purchased from Lou is only five more years.


Required:
1. Prepare the entries necessary for years 2019 through 2021 to reflect the above information.
2. Prepare a schedule showing the intangible asset section of Janes’s December 31, 2021, balance sheet.

Prepare a schedule showing the intangible asset section of Janes’s December 31, 2021, balance sheet.

Partial Balance Sheet
December 31, 2021
Intangible assets:
Total intangibles $0
Answer #1
Journal entries
Date Particulars Debit Credit
2019 patent 1050000
cash 1050000
2019 amortization expenses (1050000/10) 105000
patent 105000
2020 amortization expenses (1050000/10) 105000
patent 105000
2021 franchise 570000
cash 570000
2021 Research & development expenses 401000
cash 401000
2018 amortization exp 168000
patent 168000
2018 amortization exp (570000/10) 57000
franchise 57000
working
amortization exp-patent = 105000
for 2 years = 105000*2 = 210000
unamortized cost = (1050000-210000) = 840000
amortization exp-patent = 840000/5 = 168000
Balance sheet
intengible assets
patent 1050000
less: amortization -378000
patent -net 672000
Franchise 570000
less: amortization -57000
franchise-net 513000
total intengible 1185000